OpenStudy (anonymous):

Y=x^2-18x-59 Convert into vertex form. Thanks

OpenStudy (anonymous):

I need to do it step by step

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Im struggling with the formula

OpenStudy (jdoe0001):

is just a matter of "completing the square" for a "perfect square trinomial" I'd think you've covered this but anyhow a perfect square trinomial looks like -> \(\large x^2\pm2xy+y^2 \implies (x\pm y)^2 \)

OpenStudy (anonymous):

Y=(x^2-18x )-59 Y=(x^2-18x+81)-59 y=X^2 -18X+81-81)-59

OpenStudy (jdoe0001):

that's correct :)

OpenStudy (jdoe0001):

so you have \( (x-9)^2 -81-59\)

OpenStudy (anonymous):

I get lost at this point

OpenStudy (anonymous):

thats the answer I get but the back of the book says the answer is x^2+6x=9

OpenStudy (jdoe0001):

Y=(x^2-18x )-59 Y=(x^2-18x+81)-59 y=X^2 -18X+81-81)-59 so you have $$ y=(x^2-18x+81)-81-59\\ y=(\color{red}{x}^2-2(\color{blue}{9})(\color{red}{x})+\color{blue}{9}^2)-81-59\\ y=(\color{red}{x}-\color{blue}{9})^2-81-59 $$

OpenStudy (anonymous):

sorry x^2-4x+4

OpenStudy (jdoe0001):

hmm, that doesn't like it'd be it :/

OpenStudy (jdoe0001):

what the above gives is (x-9)^2 -140

OpenStudy (jdoe0001):

you can always recheck that in your calculator, plot the equation, check where the vertex (h,k) values are at

OpenStudy (anonymous):

Thanks for the help. This example help out alot

OpenStudy (jdoe0001):

yw